Solve the following system of equations using elimination. 5x+y=24-5x -4y=-56
Answer by Boreal(15235)

It's not clear as the equations are written without a space between them.
If it is this
5x+y=24
-5x -4y=-56
add them
x disappears
-3y=-32
y=32/3
then 5x+32/3=72/3
5x=40/3
x=8/3
If the equations are different, please resubmit